A current of 3 A flows for 50 h through a 6 Ω resistor. The energy consumed by the resistor is


1
Expert's answer
2021-08-24T14:47:28-0400
"E=I^2Rt\\E=(3)^2(6)(50\cdot3600)=9.72\ MJ"